ABC is a triangle. AB=6cm and BC=4cm. The angle ABC between these two sides is 130 degree. the missing side length AC of the triangle is?
1
Expert's answer
2012-02-16T07:27:38-0500
Let's use the law of cosines:

AC² = AB² + BC² - AB*BC*cos(ABC) = 6² + 4² - 6*4*cos(130) = 52+24*0.6428 = 67.4269cm².

So, AC = sqrt(67.4269) = 8.2114cm.